OpenStudy (anonymous):

Gardeners expect eighty-four percent of planted seeds to germinate. What fraction of seeds do gardeners expect to germinate? a. 19/15 b. 18/25 c. 21/25 d. 20/25

OpenStudy (anonymous):

first change the percentage into a fraction then simplify

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

84%= 0.84 = 84/100

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

Just simplify all the fractions into decimals: a.1.27 b.0.72 c.0.84 d.0.80 Now find which one is 84% equivalent to.

OpenStudy (anonymous):

same thing but backwards?? teach her the right way

OpenStudy (anonymous):

0.84?

OpenStudy (anonymous):

84/100 = 21/25 = c

OpenStudy (anonymous):

Yup